OpenStudy (anonymous):

Solve: 2a/3 + 1/2 = (6a-1)/6 I hit a mental block involving equations with fractions...somebody care to explain how to do this one?

OpenStudy (gw2011):

Multiply each item on both sides of the equation by 6 in order to eliminate the denominators which gives you the following: (2a)(2) + (1)(3) = 6a - 1 4a + 3 = 6a - 1 3 + 1 = 6a - 4a This is combining like terms 4 = 2a Rearranging this equation you get: 2a = 4 a = 2